# ../../Bibliographies/main.bib

### v1

@@ 4.1. Ordinal Terms
    • $\beta<_{\kappa}\omega^{\alpha}$ if $\beta<\alpha$,
    • $\omega^{\alpha}<_{\kappa}\beta$ if $\alpha\leq\beta$,
  • $\Omega^{(J)}<_{\kappa}\Omega^{(J^{\prime})}$ if $J^{\prime}<J$,
  • $\Xi^{(J)}(\alpha)<_{\kappa}\Omega^{(J^{\prime})}$ if $J^{\prime}<J$ or $J^{\prime}=J$ and $\kappa=\Omega$,
  • $\Omega^{(J)}<_{\kappa}\Xi^{(J^{\prime})}(\alpha)$ if $J^{\prime}<J$ or $J^{\prime}=J$ and $\kappa=\Xi$,
  • $\Xi^{(J)}(\alpha)<_{\kappa}\Xi^{(J^{\prime})}(\beta)$ if $J^{\prime}<J$ or $J=J^{\prime}$ and $\alpha<_{\Omega}\beta$,
  • $v^{(J)}<_{\kappa}\Omega^{(J^{\prime})}$ if $J^{\prime}<J$ or $J^{\prime}=J$ and $\kappa=\Omega$,
  • $v^{(J)}<_{\kappa}\Xi^{(J^{\prime})}(\alpha)$ if $J\leq J^{\prime}$,
  • if $\beta$ has the form $\Omega^{(J)}$, $\Xi^{(J)}(\beta^{\prime})$, or $v^{(J)}$, then $\beta<_{\kappa}\vartheta_{\Omega}\alpha$ if there is a $\gamma\in K^{\geq 0,\Xi}\alpha$ with $\beta\leq_{\kappa}\gamma$,
  • $\vartheta_{\Omega}\alpha<_{\kappa}\vartheta_{\Omega}\beta$ if:
    • $\alpha<_{\Omega}\beta$ and, for all $\gamma\in K^{\geq 0,\Xi}\alpha$, $\gamma<_{\kappa}\vartheta_{\Omega}\beta$, or
    • $\beta<_{\Omega}\alpha$ and there is some $\gamma\in K^{\geq 0,\Xi}\beta$ so that $\vartheta_{\Omega}\alpha\leq_{\kappa}\gamma$,
  • $\vartheta_{\Xi}\alpha<_{\Omega}\vartheta_{\Omega}\beta$ if there is a $\gamma\in K^{\geq 0,\Xi}\beta$ with $\vartheta_{\Xi}\alpha\leq_{\Omega}\gamma$,
  • $\vartheta_{\Omega}\alpha<_{\Omega}\vartheta_{\Xi}\beta$ if for all $\gamma\in K^{\geq 0,\Xi}\alpha$ we have $\gamma<_{\Omega}\vartheta_{\Xi}\alpha$,
  • $\vartheta_{\Xi}\alpha<_{\Xi}\vartheta_{\Omega}\beta$ if for all $\gamma\in K^{\geq 0,\Omega}\beta$ we have $\gamma<_{\Xi}\vartheta_{\Omega}\beta$,
  • $\vartheta_{\Omega}\alpha<_{\Xi}\vartheta_{\Xi}\beta$ if there is a $\gamma\in K^{\geq 0,\Omega}\beta$ with $\vartheta_{\Omega}\alpha\leq_{\Xi}\gamma$,
  • if $\beta$ has the form $\Omega^{(J)}$, $\Xi^{(J)}(\beta^{\prime})$, or $v^{(J)}$, then $\beta<_{\kappa}\vartheta_{\Xi}\alpha$ if there is a $\gamma\in K^{\geq 0,\Omega}\alpha$ with $\beta\leq_{\kappa}\gamma$,
  • $\vartheta_{\Xi}\alpha<_{\kappa}\vartheta_{\Xi}\beta$ if:
@@
    • there is a decomposition $\alpha=(\vartheta_{\Xi}\alpha^{\prime})[v_{i}\mapsto^{0,\kappa}\alpha_{i}]$ so that each $\alpha_{i}<_{\kappa}\vartheta_{\Xi}\beta$ and there is some $\gamma\in K^{\geq 0,\Xi}\beta$ so that $\vartheta_{\Xi}\alpha^{\prime}\leq_{\kappa}\gamma$, or
    • $\alpha<_{\Xi}\beta$ and for any decomposition $\beta=(\vartheta_{\Xi}\beta^{\prime})[v_{i}\mapsto^{0,\kappa}\beta_{i}]$ where each $\beta_{i}<_{\kappa}\vartheta_{\Xi}\alpha$ and any $\gamma\in K^{\geq 0,\Xi}\alpha$, we have $\vartheta_{\Xi}\beta^{\prime}\not\leq_{\kappa}\gamma$.
Because the variables are not comparable to everything, clearly this is not a linear order. The closed ordinal terms (those without variables) are linearly ordered, as can be checked by a straightforward induction.
More generally, we have the following.
@@
If $\beta$ is closed then, for any $\alpha$, there is some decomposition $\beta=\beta^{\prime}[v_{i}\mapsto\beta_{i}]$ so that $\alpha$ and $\beta^{\prime}$ are comparable.
Proof.
By simultaneous induction on the construction of $\alpha$ and $\beta$. The main case is $\vartheta_{\Xi}\alpha$ and $\vartheta_{\Xi}\beta$.
First, if there is any way to decompose $\vartheta_{\Xi}\beta$ so that the first condition in the definition verifies $\vartheta_{\Xi}\alpha<_{\kappa}\vartheta_{\Xi}\beta^{\prime}$, we may take it. That is, if there is any way to write $\vartheta_{\Xi}\beta=(\vartheta_{\Xi}\beta^{\prime})[v_{i}\mapsto^{0,\kappa}\beta_{i}]$ and $\vartheta_{\Xi}\alpha=(\vartheta_{\Xi}\alpha^{\prime})[v_{i}\mapsto^{0,\kappa}\alpha_{i}]$ so that each $\alpha_{i}<_{\kappa}\vartheta_{\Xi}\beta^{\prime}$ and there is some $\gamma\in K^{\geq 0,\Xi}\beta$ so that $\vartheta_{\Xi}\alpha^{\prime}\leq_{\kappa}\gamma$.
So we assume this is not the case. Suppose there is any decomposition $\vartheta_{\Xi}\beta=(\vartheta_{\Xi}\beta^{\prime})[v_{i}\mapsto^{0,\kappa}\beta_{i}]$ and any $\gamma\in K^{\geq 0,\Xi}\alpha$ with $\vartheta_{\Xi}\beta^{\prime}\leq_{\kappa}\gamma$. By inductively decomposing the $\beta_{i}$, we may make them comparable to $\vartheta_{\Xi}\alpha$; if we have $\vartheta_{\Xi}\alpha\leq_{\kappa}\beta^{\prime}_{i}$ for any of these decompositions then we would be in the first case, so we must have $\beta^{\prime}_{i}<_{\kappa}\vartheta_{\Xi}\alpha$.
So we have $\vartheta_{\Xi}\beta=(\vartheta_{\Xi}\beta^{*})[w_{i}\mapsto^{0,\kappa}\gamma_{i}]=\left((\vartheta_{\Xi}\beta^{\prime})[v_{i}\mapsto^{0,\kappa}\beta^{\prime}_{i}]\right)[w_{i}\mapsto^{0,\kappa}\gamma_{i}]$ where each $\beta^{\prime}_{i}<_{\kappa}\vartheta_{\Xi}\alpha$ and $\vartheta_{\Xi}\beta^{\prime}\leq_{\kappa}\gamma$. Therefore $\vartheta_{\Xi}\beta^{*}<_{\kappa}\vartheta_{\Xi}\alpha$ and we are done.
So we may assume this is not the case either. Then we may choose some decomposition $\beta=\beta^{\prime}[v_{i}\mapsto^{1,\Xi}\beta_{i}]$ so that $\beta^{\prime}$ is comparable to $\alpha$. if $\alpha<_{\kappa}\beta^{\prime}$ then we have $\vartheta_{\Xi}\alpha<_{\kappa}\vartheta_{\Xi}\beta^{\prime}$ (if there were a further decomposition inverting this, we would be in the second case). If $\beta^{\prime}<_{\kappa}\alpha$ then we have $\vartheta_{\Xi}\beta^{\prime}<\vartheta_{\Xi}\alpha$ (because if there were a decomposition of $\vartheta_{\Xi}\alpha$ inverting this then we would have been in the first case). Finally, if $\beta^{\prime}=\alpha$ then of course $\vartheta_{\Xi}\beta^{\prime}=\vartheta_{\Xi}\alpha$. ∎
@@ 4.2. Variables
Finally, of course, we need to add additional variables and verify the key lemma. We need two versions of this, one for $\Omega$ and one for $\Xi$.
The $\Omega$ version is essentially the one we did before. The main thing to note is that we need to add new variables for it, since we need variables analogous to $\Omega^{(J)}$ terms.
We add new $\Omega$-variables $v^{(J)}_{\Omega}$ and define substitution and substitutability as in Section 3.3.
Lemma 4.10 (Key Lemma for $\Omega$).
  1. (1)
    If $\alpha<\beta$, $v$ is $0,\Xi$-substitutable in $\alpha$ and $\beta$, and $\overline{\mathbb{FC}}(\gamma)<_{\Xi}(0,\Omega)$ then $\alpha[v\mapsto\gamma]<_{\Xi}\beta[v\mapsto\gamma]$.
  2. (2)
    If $\alpha<_{\Xi}\beta$ and $K^{\geq 0,\Omega}_{\Omega}\alpha<_{\Xi}\vartheta_{\Omega}\beta$ then $\vartheta_{\Omega}\alpha<_{\Xi}\vartheta_{\Omega}\beta$.
  3. (3)
    If $\alpha<_{\Xi}\beta$, $\gamma<_{\Xi}\beta$, $v$ is $0,\Xi$-substitutable in $\beta$, $K^{\geq 0,\Omega}_{\Omega}\alpha<_{\Xi}\vartheta_{\Omega}\beta$, and $K^{\geq 0,\Omega}_{\Omega}\gamma<_{\Xi}\vartheta_{\Omega}\beta$, then $\vartheta_{\Omega}(\gamma[v\mapsto\vartheta_{\Omega}\alpha])<_{\Xi}\vartheta_{\Omega}\beta$.
The proofs are just as before.
The $\Xi$ version requires a crucial novelty: we add function variables, which we call $\Xi$-variables, and we substitute terms with distinguished ordinal variables—that is, terms we view as functions—for them.
We add new terms $v_{\Xi}^{(J)}(\alpha)$, where $\alpha$ is an ordinal term.
@@
When $\alpha,\beta$ are ordinal terms, $v_{\Xi}$ is a $\Xi$-variable, and $w$ is a variable, we define $\alpha[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]$ by induction on $\alpha$:
  • #$\#\{\alpha_{i}\}[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]=\#\{\alpha_{i}[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]\}$,
  • $\omega^{\alpha}[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]=\omega^{\alpha[v\mapsto^{J,\kappa}\beta]}$,
  • $\Omega^{(J^{\prime})}[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]=\Omega^{(J^{\prime\prime})}$,
  • $\Xi^{(J^{\prime})}(\alpha)[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]=\Xi^{(J^{\prime})}(\alpha[v_{\Xi}\mapsto^{J-J^{\prime},\Omega}\beta(w)])$,
  • $(\vartheta_{\Omega}\alpha)[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]=\vartheta_{\Omega}(\alpha[v_{\Xi}\mapsto^{J,\kappa+\Omega},w]\beta)$,
  • $(\vartheta_{\Xi}\alpha)[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]=\vartheta_{\Xi}(\alpha[v_{\Xi}\mapsto^{J,\kappa+\Xi},w]\beta)$,
  • $u^{(J^{\prime})}[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]=u^{(J^{\prime})}$,
  • $(u_{\Xi}^{(J^{\prime})}(\gamma))[v_{\Xi}\mapsto^{J,\kappa}\beta(w)]=\left\{\begin{array}[]{ll}(\beta[w\mapsto^{\geq 0,\Omega}\gamma[v_{\Xi}\mapsto^{(J-J^{\prime},\Omega)}\beta(w)]])^{\geq 0,\Omega}_{+J,\kappa}&\text{if }u_{\Xi}=v_{\Xi}\\ u_{\Xi}^{(J^{\prime})}(\gamma[v_{\Xi}\mapsto^{J-J^{\prime},\Omega}\beta(w)])&\text{otherwise}\end{array}\right.$.
We define substitutability as usual.
The crucial feature of this system is that we can prove the following version of the Key Lemma, in which we substitute a function variable instead of an ordinal variable.
@@
    If $\alpha<_{\Omega}\beta$, and $v$ is $0,\Omega$-substitutable in $\gamma$ then $\gamma[v\mapsto\alpha]<_{\Omega}\gamma[v\mapsto\beta]$.
@@
    If $\alpha<_{\Omega}\beta$, $v_{\Xi}$ is $0,\Omega$-substitutable in $\alpha$ and $\beta$, $\overline{\mathbb{FC}}(\gamma)<(0,\Omega)$, and $w$ is $0,\Omega$-substitutable in $\gamma$ then $\alpha[v_{\Xi}\mapsto^{0,\Omega}\gamma(w)]<_{\Omega}\beta[v_{\Xi}\mapsto^{0,\Omega}\gamma(w)]$.
@@
    If $\alpha<_{\Omega}\beta$ and $K^{\geq 0,\Omega}_{\Xi}\alpha<_{\Omega}\vartheta_{\Xi}\beta$ then $\vartheta_{\Xi}\alpha<_{\Omega}\vartheta_{\Xi}\beta$.
@@
    If $\alpha<_{\Omega}\beta$, $\gamma<_{\Omega}\beta$, $v_{\Xi}$ is $0,\Omega$-substitutable in $\gamma$, $w$ is $0,\Xi$-substitutable in $\alpha$, $K^{\geq 0,\Omega}_{\Xi}\alpha<_{\Omega}\vartheta_{\Xi}\beta$, and $K^{\geq 0,\Omega}\gamma<_{\Omega}\vartheta_{\Xi}\beta$ then $\vartheta_{\Xi}(\gamma[v_{\Xi}\mapsto\vartheta_{\Xi}\alpha(w)])<_{\Omega}\vartheta_{\Xi}\beta$.

### v2

@@ 4.1. Ordinal Terms
    • $\beta<\omega^{\alpha}$ if $\beta<\alpha$,
    • $\omega^{\alpha}<\beta$ if $\alpha\leq\beta$,
  • $\Xi^{(J)}(\alpha)<\Xi^{(J^{\prime})}(\beta)$ if $J<J^{\prime}$ or $J=J^{\prime}$ and $\alpha<\beta$,
  • $\Xi^{(J)}(\alpha)<\vartheta\beta$ if there is a $\gamma\in K^{<0}\alpha$ with $\Xi^{(J)}(\alpha)\leq\gamma[v\mapsto^{0}0]$,
  • $\vartheta\alpha<\vartheta\beta$ if:
    • $\alpha<\beta$ and for any parameter $\beta^{\prime}$ of $\beta$ and any $\gamma\in K^{<0}\alpha$, $\gamma[v\mapsto^{0}\beta^{\prime}]<\vartheta\beta$,
@@
    • $\beta<\alpha$ and there is a parameter $\alpha^{\prime}$ of $\alpha$ and a $\gamma\in K^{<0}\beta$ so that $\vartheta\alpha\leq\gamma[v\mapsto^{0}\alpha^{\prime}]$.
One might worry about the inductive definition involved in comparisons like $\gamma[v\mapsto^{0}\beta^{\prime}]<\vartheta\beta$. The following perspective illustrates why this is justified, as well as why it yields the strong fixed point property of $\vartheta\beta$ that we need.
Evaluate the inductive definition of $<$ between $\gamma$, with the variable, and $\vartheta\beta$. This is well-defined unless we need to compare a variable in $\gamma$ with a strongly critical subterm of $\vartheta\beta$. However, since could substitute $v$ with any strongly critical subterm of $\vartheta\beta$, so we may assume $v$ is the largest strongly critical subterm of $\vartheta\beta$. In this way, we can complete the inductive comparison of $\gamma$ to $\vartheta\beta$.
Similarly, if $\gamma[v\mapsto^{0}\beta^{\prime}]<\vartheta\beta$ for all strongly critical subterms of $\vartheta\beta$, it follows that $\gamma[v\mapsto^{0}\delta]<\vartheta\beta$ for all $\delta<\vartheta\beta$, since replacing $v$ with such a $\delta$ does not change the result of the comparison.
4.2. Well-Foundedness
The basic structure of the proof is similar. We will classify ordinals based on their largest cardinal term—that is, given an ordinal $\alpha$ with $\overline{\mathbb{FC}}^{\leq 0}(\alpha)>-\infty$, we will shift it over so that $\overline{\mathbb{FC}}^{\leq 0}(\alpha^{*})=0$. We will further identify the largest $\beta$ so that $\Xi^{(0)}(\beta)$ appears $\alpha^{*}$.
Definition 4.9.
The fine cardinality of $\alpha$ with $\overline{\mathbb{FC}}^{\leq 0}(\alpha)\in\{-\infty,0\}$ is $\infty$- ∞ if $\overline{\mathbb{FC}}^{\leq 0}(\alpha)=-\infty$ and is $\beta$, where $\beta$ is maximal so that $\alpha=\alpha^{\prime}[v\mapsto^{0}\Xi^{(0)}(\beta)]$ for some $\alpha^{\prime}$ with $v$ appearing in $\alpha^{\prime}$, if $\overline{\mathbb{FC}}^{\leq 0}(\alpha)=0$. We write $\kappa(\alpha)$ for the fine cardinality of $\alpha$.
Definition 4.10.
We define a sequence of sets inductively:
  • $M_{-\infty}$ is all ordinal terms with formal cardinality $\infty$- ∞,
  • given $M_{c}$, we let $\mathrm{Acc}_{c}$ be the well-founded part of $M_{c}$,
  • for any $c=(n,\beta)$ so that $\beta=\beta^{*}\in\mathrm{Acc}_{(-\mathbb{G}(\beta^{*}),\kappa(\beta^{*}))}$, we let $M_{c}$ be the set of ordinal terms $\alpha$ with $\overline{FC}^{\leq 0}(\alpha)=0$, $\mathbb{G}(\alpha)\geq-n$, $\kappa(\alpha)=\beta$, and $\{\gamma^{*}\mid\gamma\in K^{<0}\alpha\}\subseteq\bigcup_{c^{\prime}<c}\mathrm{Acc}_{c^{\prime}}$
where $-\infty<(n,\beta)$ for all $n,\beta$ and $(n,\beta)<(n^{\prime},\beta^{\prime})$ if $n<n^{\prime}$ or $n=n^{\prime}$ and $\beta<\beta^{\prime}$.
As before, we have:
@@ 4.2. Well-Foundedness
  • Whenever $c\leq c^{\prime}$, $\alpha\in\mathrm{Acc}_{c}$, $\beta\in\mathrm{Acc}_{c^{\prime}}$, we have $\alpha^{\geq 0}_{-k}\#\beta\in\mathrm{Acc}_{c^{\prime}}$ where, if $c=(m,\alpha^{\prime})$ and $c^{\prime}=(n,\beta^{\prime})$ then $k=n-m$ (and if $c$ does not have this form then $c=-\infty$ and $k$ does not matter).
  • If $\alpha\in\mathrm{Acc}_{c}$ then $\omega^{\alpha}\in\mathrm{Acc}_{c}$.
  • If $\alpha\in\mathrm{Acc}_{c}$ then $\{(\vartheta\alpha)^{*},(\vartheta(\alpha^{\leq 0}_{-1}))\}\subseteq\bigcup_{c^{\prime}}\mathrm{Acc}_{c^{\prime}}$.
@@ 4.3. Variables
Finally, of course, we wish to verify the key lemma. We already have variables, but we need additional ones which are properly like the $\Xi$ terms—that is, we need variables which are functions.
So we extend our system by terms $V^{(J)}(\alpha)$ where $\alpha$ is an ordinal term and $V$ is our new variable and $V^{(J)}(\alpha)$ is below $\Xi^{(J)}(\alpha)$. As usual, we do not allow $V$ to appear inside $\vartheta$.
@@
When $\alpha$ are ordinal terms and $\beta$ is an ordinal term with a distinguished (ordinal) variable $w$ which is 00-substitutable in $\beta$, we define $\alpha[V\mapsto^{J}\beta(w)]$ by induction on $\alpha$:
  • #$\#\{\alpha_{i}\}[V\mapsto^{J}\beta(w)]=\#\{\alpha_{i}[V\mapsto^{J}\beta(w)]\}$,
  • $\omega^{\alpha}[V\mapsto^{J}\beta(w)]=\omega^{\alpha[V\mapsto^{J}\beta(w)]}$,
  • $\Xi^{(J^{\prime})}(\alpha)[V\mapsto^{J}\beta(w)]=\Xi^{(J^{\prime})}([V\mapsto^{J-J^{\prime}}\beta(w)])$,
  • $(\vartheta\alpha)[V\mapsto^{J}\beta(w)]=\vartheta(\alpha[V\mapsto^{J-1}\beta(w)])$,
  • $V^{(J^{\prime})}(\alpha)[V\mapsto^{J}\beta(w)]=\beta(\alpha[V\mapsto^{0}\beta(w)])$.
We define when $V$ is $J$-substitutable in $\alpha$ by:
  • $V$ is $J$-substitutable in #$\#\{\alpha_{i}\}$ if $V$ is $J$-substitutable in each $\alpha_{i}$,
  • $V$ is $J$-substitutable in $\omega^{\alpha}$ if $V$ is $J$-substitutable in $\alpha$,
  • $V$ is $J$-substitutable in $\Xi^{(J^{\prime})}(\alpha)$ if $J\leq J^{\prime}$ and $V$ is $J-J^{\prime}$-substitutable in $\alpha$,
  • $V$ is $J$-substitutable in $\vartheta\alpha$ if $V$ does not appear in $\vartheta\alpha$,
  • $V$ is $J$-substitutable in $V^{(J^{\prime})}(\alpha)$ if $J=J^{\prime}$ and $V$ is 00-substitutable in $\alpha$.
Definition 4.13.
When $\overline{\mathbb{FC}}(\gamma)<0$, we define $D_{\gamma}(\beta)=D_{0,\gamma}(\beta)=\vartheta(\omega^{\Xi^{(0)}(1)\#\beta}\#\gamma(\Xi^{(0)}(0)))$ and $D_{m+1,\gamma}(\beta)=D_{0,0}(D_{m,\gamma}(\beta))$.
When $\overline{\mathbb{FC}}(\gamma)<0$, we define $\alpha\ll_{\gamma}\beta$ to hold if $\alpha<\beta$ and every $\eta\in K^{<0}\alpha$ is bounded by $D_{m,\gamma}(\beta)$ where $m$ is least so that $\overline{\mathbb{FC}}(\gamma)\leq\overline{\mathbb{FC}}(\eta)$.
We expect that $\eta$ has free variables; we understand this to be the comparison in which the free variables of $\eta$ are larger than any proper strongly critical subterm of $D_{m,\gamma}(\beta)$.
@@
    If $\alpha<\beta$, and $v$ is 00-substitutable in $\gamma$ then $\gamma[v\mapsto^{0}\alpha]<\gamma[v\mapsto^{0}\beta]$.
@@
    If $\overline{\mathbb{FC}}(\delta)<0$, $\alpha\ll_{\delta}\beta$, $V$ is 00-substitutable in $\alpha$ and $\beta$, $\overline{\mathbb{FC}}(\gamma)<0$, and $w$ is 00-substitutable in $\gamma$ then $\alpha[V\mapsto^{0}\gamma(w)]<_{\Omega}\beta[V\mapsto^{0}\gamma(w)]$.
@@
    If $\overline{\mathbb{FC}}(\delta)<0$ and $\alpha\ll_{\delta}\beta$ are ordinal terms so that every free variable is $J$-substitutable for some $J<0$ then $D_{\delta}(\alpha)\ll_{0}D_{\delta}(\beta)$.
@@
    If $\overline{\mathbb{FC}}(\delta)<0$, $\alpha\ll_{\delta}\beta$, $\gamma\ll_{\delta}\beta$, $V$ is 00-substitutable in $\gamma$, $V$ does not appear in $\beta$, $w$ is 00-substitutable in $\alpha$, then $D_{0,D_{0,\delta}(\alpha)^{\leq 0}_{-1}}(\gamma[V\mapsto^{0}D_{0,\delta}(\alpha)^{\leq 0}_{-1}])\ll_{0}D_{0,\delta}(\alpha)$.
Proof.
  1. (1)
    By induction on $\gamma$.
  2. (2)
    By induction on the comparison between $\alpha$ and $\beta$. The case which would otherwise be complicated is avoided since $V$ does not appear inside any $\vartheta$.
  3. (3)
    Since $\alpha<\beta$ we have $\omega^{\Xi^{(0)}(1)\#\alpha}\#\delta(\Xi^{(0)}(0))<\omega^{\Xi^{(0)}(1)\#\beta}\#\delta(\Xi^{(0)}(0))$. If $\eta\in K^{<0}(\omega^{\Xi^{(0)}(1)\#\alpha}\#\delta(\Xi^{(0)}(0)))$ then $\eta\in K^{<0}(\alpha)\cup\{\delta\}$ and is therefore bounded by $D_{\gamma}(\beta)$, and furthermore by $D_{m,\gamma}(\beta)$ for the suitable $m$.
  4. (4)
    We have $\omega^{\Xi^{(0)}(1)\#\gamma[V\mapsto^{0}D_{0,\delta}(\alpha)^{\leq 0}_{-1}]}\#D_{0,\delta}(\alpha)^{\leq 0}_{-1}(\Xi^{(0)}(0))<\omega^{\Xi^{(0)}(1)\#\beta}\#D_{0,\delta}(\alpha)^{\leq 0}_{-1}(\Xi^{(0)}(0))$ (using (1) to show that $\gamma[V\mapsto^{0}D_{0,\delta}(\alpha)^{\leq 0}_{-1}]<\beta$). Suppose $\eta\in K^{<0}(\omega^{\Xi^{(0)}(1)\#\gamma[V\mapsto^{0}D_{0,\delta}(\alpha)^{\leq 0}_{-1}]}\#D_{0,\delta}(\alpha)^{\leq 0}_{-1}(\Xi^{(0)}(0)))$. Then $\eta\in K^{<0}\gamma\cup\{D_{0,\delta}(\alpha)^{\leq 0}_{-1}(w)\}$ and is therefore bounded by the suitable $D_{m,\gamma}(\beta)$.